OpenStudy (anonymous):

I need help with this problem Graph y-2=-1/3(x+3)^2 and find the following I have already made my graph and found the vertex, y-intercept, Max value, axis of symmetry. But I cant figure out how to calculate the x-intercept.

OpenStudy (anonymous):

Vertex:-3,2 Max:2 Axis of Symmetry: -3 y-intercept: (0,-1)

OpenStudy (phi):

the x-intercept means y = 0 so we can replace y with 0 in your equation to start with: \[ y-2=-\frac{1}{3} (x+3)^2 \\ 0-2=-\frac{1}{3} (x+3)^2 \\ -2=-\frac{1}{3} (x+3)^2\] I would multiply both sides by -3 to get \[ -3 \cdot -2=-3 \cdot -\frac{1}{3}\cdot (x+3)^2 \\ 6= (x+3)^2 \] or \[ (x+3)^2 = 6\] I would take the square root of both sides as the next step.

OpenStudy (anonymous):

Thank you, phi!